OpenStudy (anonymous):

Use a table of values to graph each equation. State domain and range 2x^2+4x-6

OpenStudy (anonymous):

See the youtube video. http://www.youtube.com/watch?v=c74m88mbi9k The domain is \[(-\infty,+\infty)\] The range is \[[-12,+\infty)\]

jhonyy9 (jhonyy9):

2x^2 +4x -6=0 /2 x^2 +2x -3=0 (x+3)(x-1)=0 x+3=0 x=-3 x-1=0 x=1